# On the choice of the non-trainable internal weights in random feature maps

### v1

@@ 4. How to sample good internal weights
$L_{0}<|\mathbf{W}_{\rm in}\mathbf{u}_{n}+\mathbf{b}_{\rm in}|<L_{1},\qquad\forall\;n=1,2,\ldots N.$ (12)
@@
$L_{0}<|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|<L_{1},\qquad\forall\;n=1,2,\ldots N.$ (13)
@@
$|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|\leq L_{0},\qquad\forall\;n=1,2,\ldots N,$ (14)
@@
$|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|\geq L_{1},\qquad\forall\;n=1,2,\ldots N.$ (15)
@@
This categorization of rows of the internal parameters is useful for investigating the effects of different realizations of the random feature map on its forecasting skill. Note that this is not an exhaustive classification since there exist rows that satisfy different inequalities for different observations ${\mathbf{u}}_{n}$ and do not satisfy (13) for the whole data set $\mathbf{U}$. Although not considered here, such mixed rows may be an interesting topic for further exploration.
@@
$S_{-}\,$ =$=\,\{(\mathbf{w},b)\in\mathbb{R}^{D+1}:-L_{1}<\mathbf{w}\cdot\mathbf{u}_{n}+b<-L_{0}\;\;\forall\;n=1,2,\ldots,N\},$ (17)
$S_{+}\,$ =$=\,\{(\mathbf{w},b)\in\mathbb{R}^{D+1}:+L_{0}<\mathbf{w}\cdot\mathbf{u}_{n}+b<+L_{1}\;\;\forall\;n=1,2,\ldots,N\}.$ (18)
@@
it suffices to sample from only one of these convex sets and then uniformly sample the sign of the internal weights to sample from $\Omega_{g}$. Hence, the sampling problem is effectively a convex problem. Analogously, we define $\Omega_{l}$ and $\Omega_{s}$ to be the solution sets to the problems (14) and (15) respectively, and again sampling these sets are also convex problems.
We present in the next two subsections algorithms to effectively sample from the sets $\Omega_{g,l,s}$. A naive choice of sampling algorithm would be to uniformly sample from the $D$-dimensional hypercube with the $2^{D}$ corners defined by the observed extremal training data points, and checking the inequality (12), if we want to sample from $\Omega_{g}$, let’s say. This, however, is computationally very costly as typically the solution set only occupies a small region within that hypercube. Instead, we begin with a hit-and-run algorithm sampling from $\Omega_{g}$ in Section 4.1 and then present a faster more efficient hit-and-run algorithm to sample from an equivalent restricted solution set in Section 4.2.
@@ 4.1. Standard hit-and-run sampling of good internal parameters
We now describe a computationally cheap and easy to implement numerical algorithm to uniformly sample from the solution sets $\Omega_{g,l,s}$. We shall employ hit-and-run algorithms [40, 43]. To uniformly sample a set $\Omega$ with hit-and-run, one starts from a feasible point inside the set, considers the line through that point in a randomly chosen direction, and then randomly picks a point on the intersection of that line and the set $\Omega$ as a new point. This process is then repeated to generate further samples. For convex sets the hit-and-run samples converge to uniform samples in total variation distance. The convergence depends polynomially on the number of iterations and dimension with the polynomial dependence on dimension being of low order [26, 2, 27]. This and the fast mixing properties make hit-and-run algorithms an attractive method to uniformly sample from $\Omega_{g,l,s}$.
We sample the augmented internal weight matrix $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ row by row. Each sample lies then in a $D+1$-dimensional search space for $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$. Due to (19) it suffices to sample from $S_{+}$. In order to perform hit-and-run, given a point, we need to efficiently determine if a point lies in $S_{+}$. Focusing on a convex conical subset of $S_{+}$, it turns out that we can determine if a point belongs to $S_{+}$ by checking just two inequalities. Define the convex cone
@@
To initialize the hit-and-run algorithm with a feasible point we choose $(\mathbf{w},b)=(\mathbf{0},b_{0})\in S_{+}$ for $b_{0}\in(L_{0},L_{1})$. To determine the line segments inside $S_{+}$ we use bisection together with the feasibility criterion (23). The hit-and-run algorithm requires a few iterations to ensure that the samples become independent of the initial feasible weight point $(\mathbf{w},b)=(\mathbf{0},b_{0})$.
@@
1:  Input: training data $\mathbf{U}$.
2:  Choose number of decorrelation iterations $K\in\mathbb{N}$ and $L_{0},L_{1}\in\mathbb{R}_{>0}$. Below $\pi$ denotes the canonical projection: $\pi(\mathbf{w},b)=\mathbf{w}$.
@@
8:     $A\leftarrow\{a\in\mathbb{R}:(\mathbf{w}+a\pi(\mathbf{d}))\cdot\mathbf{x}_{-}(\operatorname{sgn}(\mathbf{w}+a\pi(\mathbf{d})))+b+a\mathbf{d}_{D+1}>L_{0}\land(\mathbf{w}+a\pi(\mathbf{d}))\cdot\mathbf{x}_{+}(\operatorname{sgn}(\mathbf{w}+a\pi(\mathbf{d})))+b+a\mathbf{d}_{D+1}<L_{1}\}$.
9:     $a_{0}\leftarrow\inf A$.
10:     $a_{1}\leftarrow\sup A$.
11:     Sample $a$ uniformly from $(a_{0},a_{1})$.
12:     $(\mathbf{w},b)\leftarrow(\mathbf{w},b)+a\mathbf{d}$.
13:     $k\leftarrow k+1$.
14:  end while
15:  Uniformly sample a scalar $z$ from $\{-1,1\}$ to determine which set to sample from, $S_{-}$ or $S_{+}$.
16:  if $z=1$ then
17:     $(\mathbf{w},b)$ is our final good row sample.
18:  else
19:     $(-\mathbf{w},-b)$ is our final good row sample.
20:  end if
@@ 4.2. One-shot hit-and-run sampling
To generate good (or linear or saturated) random feature maps one can restrict the solution spaces $\Omega_{g,l,s}$ by first sampling $b$ appropriately and then sampling $\mathbf{w}$ on a $D$-dimensional search space. For ease of presentation we describe the algorithm for sampling from the good set $\Omega_{g}$. We sample $b$ uniformly from the interval $(L_{0},L_{1})$. The weights $\mathbf{w}$ are then sampled from the restricted solution set $\Omega_{g}^{R}=S_{+}^{R}\cup S_{-}^{R}$ with
@@
Since $S_{-}^{R}=-S_{+}^{R}$, sampling from the nonconvex set $\Omega_{g}^{R}$ can again be done by sampling from the convex set $S_{+}^{R}$ and then multiplying the sample with $1$ or $-1$ uniformly randomly. This restriction allows us to perform hit-and-run sampling on a $D$-dimensional random convex set instead of a $(D+1)$-dimensional convex set. Note that fixing $b$ is akin to shrinking the search space from $S_{+}$ to $\pi S_{+}^{R}$ where $\pi$ is the canonical projection with $\pi(\mathbf{w},b)=\mathbf{w}$. Note that we can partition $\pi S_{+}^{R}$ according to
@@
Figure 4. Schematic of the one-shot hit-and-run algorithm 2. The weight point $\mathbf{0}$ is always an interior point of $\pi S_{+}^{R}$ and the cone $V(\mathbf{s})$ is a $D$-dimensional orthant. The set $\pi S_{+}^{R}$ is drawn as bounded here, but it can be unbounded depending on the training data $\mathbf{u}_{n}$.
@@
Algorithm 2 summarizes the one-shot hit-and-run sampling of a good row described above. Note that, depending on the training data $\mathbf{U}$ it is possible for $\pi S_{+}^{R}$ to be unbounded which is why +$\infty$+ ∞ appears in the algorithm. We can extend the notion of restriction to the coordinates of $\mathbf{w}$ as well by restricting the intervals where we are allowed to sample them from which is akin to regularizing parameters in machine learning [19, 18, 13] but we do not consider such algorithms here. Obvious modifications of Algorithm 1 and Algorithm 2 let us sample linear and saturated rows which we refrain from describing here to avoid repetition.
1:  Input: training data $\mathbf{U}$.
@@ 4.3. Performance of the hit-and-run sampling
Figure 5. Empirical histograms for samples generated using standard and one-shot hit-and-run algorithms 1 and 2, respectively. The left panel shows the distributions of the entries of $\mathbf{W}_{\rm in}$ and the right panel shows the distributions of the entries of $\mathbf{b}_{\rm in}$. For each algorithm $500$ rows of internal parameters were generated. Algorithm 1 used $K=10$ decorrelation iterations.
@@
The two hit-and-run algorithms 1 and 2 are designed to uniformly sample from the set of good rows $(\mathbf{w},b)$. The resulting distributions for the weights and biases are shown in Figure 5. For the standard hit-and-run algorithm 1 it was found that $K=10$ decorrelation steps were sufficient and results were very similar for $K=100$ iterations. It is clearly seen that the distributions are far from being the usually employed uniform or Gaussian distribution. The distributions are very similar for both algorithms. In particular, the standard Algorithm 1 exhibits the same lack of biases with small absolute value, as the one-shot hit-and-run Algorithm 2.
@@
Whereas the one-shot hit-and-run algorithm excludes biases with absolute values smaller than $L_{0}$ by design, this may seem surprising for the standard hit-and-run algorithm. This can be explained as follows. For $0<b<L_{0}$ and $(\mathbf{w},b)\in S_{+}$ we require that $\mathbf{w}\cdot\mathbf{u}$ lies in the positive interval $(L_{0}-b,L_{1}-b)$ for all training data $\mathbf{u}$. Since the directions of the vectors ${\mathbf{u}}$ in the training data are typically distributed over some range, $\mathbf{w}\cdot\mathbf{u}$ is typically not sign-definite for all data points ${\mathbf{u}}$. This implies that for all parameters in $\Omega_{g}$ we typically have $|b|>L_{0}$; a similar argument shows that typically $|b|<L_{1}$. Hence, for typical data $\mathbf{u}$ $\Omega_{g}^{R}=\Omega_{g}$ and the search space of the one-shot hit-and-run algorithm 2 is the same as that of the standard algorithm 1. We have tested that the forecasting skill is the same for both sampling algorithms.
@@
For the weights and biases which were obtained by sampling uniformly from an interval as in Figure 1, we checked that the weights corresponding to high forecasting skill indeed all satisfy our criterion of being good rows (12). This highlights the advantage of our non-parametric sampling over sampling strategies involving a set of parametrized distributions. In applications we hence use the computationally more efficient one-shot hit-and-run algorithm 2.
@@
The hit-and-run algorithms 1 and 2 were designed to uniformly sample from the sets $\Omega_{g,l,s}$. This does not imply, however, that ${\mathbf{w}}_{\rm{in}}{\mathbf{u}}+{\mathbf{b}}_{\rm{in}}$ is uniformly distributed in the interval $(-L_{1},-L_{0})\cup(L_{0},L_{1})$. To quantify the occupied range of random features we introduce the following notation. A sample $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ produces outputs the absolute values of which lie in the interval $[m_{i},M_{i}]$, i.e.
@@
Figure 6 shows that the features only occupy a relatively small part of the desired interval and the observed maximum value of the effective range $\mathcal{R}\approx 0.52$ is much smaller than the desired range of $L_{1}-L_{0}=3.1$. This reduction of the range $\mathcal{R}$ can be explained by a simple approximate model. Consider the case when the extreme values $m_{i}$ and $M_{i}$ are i.i.d random variables, and $M_{i}$ is drawn uniformly according to $M_{i}\sim\mathcal{U}[L_{0},L_{1}]$. The lower bound $m_{i}$ is then conditionally distributed according to $m_{i}\sim\mathcal{U}[L_{0},M_{i}]$.
By the central limit theorem, for $D_{r}\gg 1$ the effective range (28) is a normally distributed random variable with mean
$\mathbb{E}[\mathcal{R}]\approx$ $\,\mathbb{E}[M_{i}]-\mathbb{E}[m_{i}|M_{i}]=\,\mathbb{E}[M_{i}]-\mathbb{E}\left[\frac{L_{0}+M_{i}}{2}\right]$
$=$ $\,\frac{L_{1}-L_{0}}{4}.$ (29)
Hence the range is approximately $\mathbb{E}[\mathcal{R}]=(L_{1}-L_{0})/4$ which is significantly reduced from the desired range $L_{1}-L_{0}$. Similarly, by the central limit theorem, the standard deviation of the effective range converges to
$\sigma[\mathcal{R}]\approx\frac{\sigma[M_{i}-m_{i}]}{\sqrt{D_{r}}},$ (30)
which implies a concentration of the effective range $\mathcal{R}$ around the mean (29) for high feature dimensions $D_{r}$, consistent with the observations in Figure 6.
One would like the range $\mathcal{R}$ to be as large as possible. Indeed, Figure 7 shows that $\mathcal{R}$ and $\tau_{f}$ are positively correlated justifying our assumption that better exploration of the space of good features generally yields better forecasting skill. The observation of a small effective range $\mathcal{R}$ suggests that tuning $L_{0}$ and $L_{1}$ to make the interval $(L_{0},L_{1})$ larger may be beneficial for the forecasting skill of the random feature map.
Figure 6. Empirical histogram of the effective range $\mathcal{R}$ of random features with only good rows for $L_{0}=0.4$ and $L_{1}=3.5$. We used $500$ independent samples and a feature dimension of $D_{r}=300$. The mean is estimated as $\mathbb{E}[\mathcal{R}]=0.44$ and the standard deviation as $\sigma[\mathcal{R}]=0.03$.
@@
Figure 7. Scatter plot of the effective range $\mathcal{R}$ and the forecast time $\tau_{f}$ when only good rows are used with $p_{g}=1$. Each dot represents a different realization of the random feature map which was trained and tested on randomized data. We observe a Pearson correlation coefficient of $0.37$ between $\tau_{f}$ and $\mathcal{R}$. The feature dimension is $D_{r}=300$ and a regularization parameter of $\beta=4\times 10^{-5}$ is used with training data length $N=20,000$.
5. Results

### v2

@@ 4. How to sample good internal weights
$L_{0}<|\mathbf{W}_{\rm in}\mathbf{u}_{n}+\mathbf{b}_{\rm in}|<L_{1},\qquad\forall\;n=0,1,\ldots N.$ (12)
@@
$L_{0}<|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|<L_{1},\qquad\forall\;n=0,1,\ldots N.$ (13)
@@
$|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|\leq L_{0},\qquad\forall\;n=0,1,\ldots N,$ (14)
@@
$|\mathbf{w}_{i}^{\rm in}\cdot\mathbf{u}_{n}+b_{i}^{\rm in}|\geq L_{1},\qquad\forall\;n=0,1,\ldots N.$ (15)
@@
This categorization of rows of the internal parameters is useful for investigating the effects of different realizations of the random feature map on its forecasting skill. Note that this is not an exhaustive classification since there exist rows that satisfy different inequalities for different observations ${\mathbf{u}}_{n}$ and do not satisfy (13) for the whole data set $\mathbf{U}$. Although not considered here, such mixed rows may be an interesting topic for further exploration. The distinction into “good” (nonlinear, non-saturated) and “bad” (linear or saturated) parameters only pertains to the task of longest possible sequential forecasting. For example, for the task of predicting the state for only a single step, all parameters, good or bad, perform equally well.
@@
$S_{-}\,$ =$=\,\{(\mathbf{w},b)\in\mathbb{R}^{D+1}:-L_{1}<\mathbf{w}\cdot\mathbf{u}_{n}+b<-L_{0}\;\;\forall\;n=0,1,\ldots,N\},$ (17)
$S_{+}\,$ =$=\,\{(\mathbf{w},b)\in\mathbb{R}^{D+1}:+L_{0}<\mathbf{w}\cdot\mathbf{u}_{n}+b<+L_{1}\;\;\forall\;n=0,1,\ldots,N\}.$ (18)
@@
it suffices to sample from only one of these convex sets and then uniformly sample the sign of the internal weights to sample from $\Omega_{g}$. Hence, the sampling problem is effectively a convex problem. Analogously, we define $\Omega_{l}$ and $\Omega_{s}$ to be the solution sets to the problems (14) and (15) respectively, and again sampling these sets are convex problems. We will see that the data-informed constraint for good parameters (12) allows for sufficient variability in the features (cf. Section 4.3).
We present, in the next two subsections, algorithms to effectively sample from the good set $\Omega_{g}$ to enhance the forecasting capabilities of random feature maps, as well as from the sets $\Omega_{l}$ and $\Omega_{s}$ to illustrate their effect on random feature maps in Section 5. A naive choice of sampling algorithm would be to uniformly sample from the $D$-dimensional hypercube with the $2^{D}$ corners defined by the extremal training data points, and checking the inequality (12), if we want to sample from $\Omega_{g}$, let’s say. This, however, is computationally very costly as typically the solution set only occupies a small region within that hypercube.
Our aim is to sample good parameters. There is no a priori reason to sample uniformly from the convex set of good parameters. However, as we will see, uniform sampling lends itself to an efficient implementation. Furthermore, we stress that sampling uniformly from the parameter space does not imply uniform sampling in the feature space.
We consider a standard hit-and-run algorithm sampling from $\Omega_{g}$ in Section 4.1 and then present a faster, more efficient hit-and-run algorithm to sample from an equivalent restricted solution set in Section 4.2.
@@ 4.1. Standard hit-and-run sampling of good internal parameters
We now describe a computationally cheap and easy to implement numerical algorithm to uniformly sample from the solution sets $\Omega_{g},\Omega_{l}$ and $\Omega_{s}$. We shall employ hit-and-run algorithms [47, 50]. To uniformly sample a set $\Omega$ with hit-and-run, one starts from a feasible point inside the set, considers the line through that point in a randomly chosen direction, and then randomly picks a point on the intersection of that line and the set $\Omega$ as a new point. This process is then repeated to generate further samples. For convex sets the hit-and-run samples converge to uniform samples in total variation distance. The convergence depends polynomially on the number of iterations and dimension with the polynomial dependence on dimension being of low order [32, 2, 33]. This and the fast mixing properties make hit-and-run algorithms an attractive method to uniformly sample from $\Omega_{g},\Omega_{l}$ and $\Omega_{s}$.
We sample the augmented internal weight matrix $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ row by row. Each sample lies then in a $D+1$-dimensional search space for $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$. Due to (19) it suffices to sample from $S_{+}$. In order to perform hit-and-run, given a point, we need to efficiently determine if it lies in $S_{+}$. Focusing on a convex conical subset of $S_{+}$, it turns out that we can determine if a point belongs to $S_{+}$ by checking just two inequalities. Define the convex cone
@@
To initialize the hit-and-run algorithm with a feasible point we choose $(\mathbf{w},b)=(\mathbf{0},b_{0})\in S_{+}$ for $b_{0}\in(L_{0},L_{1})$. To determine the line segments inside $S_{+}$ we use bisection together with the feasibility criterion (23). The hit-and-run algorithm requires a few iterations to ensure that the samples become independent of the initial feasible point $(\mathbf{w},b)=(\mathbf{0},b_{0})$.
@@
1:  Input: data $\mathbf{U}$.
2:  Choose number of decorrelation iterations $K\in\mathbb{N}$ and $L_{0},L_{1}\in\mathbb{R}_{>0}$.
@@
8:     Determine $A=$ the maximal line segment passing through $(\mathbf{w},b)$ along direction $\mathbf{d}$ and contained within $S_{+}$, using the feasibility criterion (23) and the bisection method.
9:     Uniformly sample $(\mathbf{w}^{\prime},b^{\prime})$ from $A$.
10:     $(\mathbf{w},b)\leftarrow(\mathbf{w}^{\prime},b^{\prime})$
11:     $k\leftarrow k+1$.
12:  end while
13:  Uniformly sample a scalar $z$ from $\{-1,1\}$ to determine which set to sample from, $S_{-}$ or $S_{+}$.
14:  if $z=1$ then
15:     $(\mathbf{w},b)$ is our final good row sample.
16:  else
17:     $(-\mathbf{w},-b)$ is our final good row sample.
18:  end if
@@ 4.2. One-shot hit-and-run sampling
To generate good (or linear or saturated) random feature maps one can restrict the solution set $\Omega_{g}$ (or $\Omega_{l}$ or $\Omega_{s}$) by first sampling $b$ appropriately and then sampling $\mathbf{w}$ from a $D$-dimensional search space. For ease of presentation, we describe the algorithm for sampling from the good set $\Omega_{g}$. We sample $b$ uniformly from the interval $(L_{0},L_{1})$. The weights $\mathbf{w}$ are then sampled from the restricted solution set $\Omega_{g}^{R}=S_{+}^{R}\cup S_{-}^{R}$ with
@@
Since $S_{-}^{R}=-S_{+}^{R}$, sampling from the nonconvex set $\Omega_{g}^{R}$ can again be done by sampling from the convex set $S_{+}^{R}$ and then multiplying the sample with $1$ or $-1$ uniformly randomly. This restriction allows us to perform hit-and-run sampling on a $D$-dimensional random convex set instead of a $(D+1)$-dimensional convex set. Note that fixing $b$ is akin to shrinking the search space from $S_{+}$ to $\pi S_{+}^{R}$ where $\pi$ is the canonical projection: $\pi(\mathbf{w},b)=\mathbf{w}$. Also note that we can partition $\pi S_{+}^{R}$ according to
@@
Figure 5. Schematic of the one-shot hit-and-run Algorithm 2. The weight point $\mathbf{0}$ is always an interior point of $\pi S_{+}^{R}$ and the cone $V(\mathbf{s})$ is a $D$-dimensional orthant. The set $\pi S_{+}^{R}$ is drawn as bounded here, but it may be unbounded depending on the training data $\mathbf{u}_{n}$.
@@
Algorithm 2 summarizes the one-shot hit-and-run sampling of a good row. Note that, depending on the training data $\mathbf{U}$, it is possible for $\pi S_{+}^{R}$ to be unbounded, which is why +$\infty$+ ∞ appears in the algorithm. We can extend the notion of restriction to the coordinates of $\mathbf{w}$ as well by restricting the intervals where we are allowed to sample them from, which is akin to regularizing parameters in machine learning [24, 23, 14] but we do not consider such algorithms here. Obvious modifications of Algorithm 1 and Algorithm 2 let us sample linear and saturated rows which we refrain from describing here to avoid repetition.
1:  Input: data $\mathbf{U}$.
@@ 4.3. Performance and comparison of the hit-and-run algorithms
Figure 6. Empirical histograms for samples generated using standard and one-shot hit-and-run Algorithms 1 and 2, respectively. The left panel shows the distributions of the entries of $\mathbf{W}_{\rm in}$ and the right panel shows the distributions of the entries of $\mathbf{b}_{\rm in}$. For each algorithm $10,000$ rows of internal parameters were generated. The standard hit-and-run Algorithm 1 used $K=10$ decorrelation iterations.
@@
The two hit-and-run Algorithms 1 and 2 are designed to uniformly sample from the $D+1$-dimensional set of good rows $\Omega_{g}$. This does not imply that the individual entries of $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ are uniformly distributed. The distributions for the entries of weights and biases sampled from the two hit-and-run algorithms are shown in Figure 6. For the standard hit-and-run Algorithm 1 it was found that $K=10$ decorrelation steps were sufficient and results were very similar for $K=100$ iterations. It is clearly seen that the distributions are far from being the usually employed uniform or Gaussian distribution. The distributions are very similar for both algorithms. In particular, the standard Algorithm 1 exhibits the same lack of biases $b$ with small absolute value, as the one-shot hit-and-run Algorithm 2.
@@
Whereas the one-shot hit-and-run algorithm excludes biases $b$ with absolute values smaller than $L_{0}$ by design, this may seem surprising for the standard hit-and-run algorithm. This can be explained as follows. For $0<b<L_{0}$ and $(\mathbf{w},b)\in S_{+}$ we require that $\mathbf{w}\cdot\mathbf{u}$ lies in the positive interval $(L_{0}-b,L_{1}-b)$ for all training data $\mathbf{u}$. Since the directions of the vectors ${\mathbf{u}}$ in the training data are typically distributed over some range, $\mathbf{w}\cdot\mathbf{u}$ is typically not sign-definite for all data points ${\mathbf{u}}$. This implies that for all parameters in $\Omega_{g}$ we typically have $|b|>L_{0}$; a similar argument shows that typically $|b|<L_{1}$. Hence, for typical data $\mathbf{u}$, we have $\Omega_{g}^{R}=\Omega_{g}$ and the search space of the one-shot hit-and-run Algorithm 2 is the same as that of the standard Algorithm 1.
@@
For the weights and biases which were obtained by sampling uniformly from an interval as in Figure 1, we checked that the weights corresponding to high forecasting skill indeed all satisfy our criterion of being good rows (12). This highlights the advantage of our non-parametric sampling over sampling strategies involving a set of parametrized distributions.
@@
The hit-and-run Algorithms 1 and 2 were designed to uniformly sample from the set $\Omega_{g}$. This does not imply, however, that ${\mathbf{w}}_{\rm{in}}{\mathbf{u}}+{\mathbf{b}}_{\rm{in}}$ is uniformly distributed in the interval $(-L_{1},-L_{0})\cup(L_{0},L_{1})$. To quantify the occupied range of random features we introduce the following notation. A sample $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ produces outputs the absolute values of which lie in the interval $[m_{i},M_{i}]$, i.e.
@@
By the central limit theorem, for $D_{r}\gg 1$ the effective range (28) approaches a normally distributed random variable. Figure 7 shows the histogram of the range $\mathcal{R}$ for the standard and the one-shot hit-and-run algorithms when only good features are used with $p_{g}=1$. As expected both algorithms generate near-Gaussian distributions of $\mathcal{R}$.
The standard hit-and-run algorithm generates a wider range with $\mathbb{E}[\mathcal{R}]=1.0$ and standard deviation $\sigma[\mathcal{R}]=0.03$ compared to the one-shot hit-and-run algorithm with $\mathbb{E}[\mathcal{R}]=0.42$ and standard deviation $\sigma[\mathcal{R}]=0.02$. One would like the range $\mathcal{R}$ to be as large as possible, assuming that higher variability in the features increases the expressivity of the random feature map, and hence the forecast skill. Figure 8 shows the histogram of the forecast times $\tau_{f}$ for the two algorithms. The standard hit-and-run algorithm exhibits a slightly better forecast skill as measured by an approximately $6\%$ larger mean forecast time of $\mathbb{E}[\tau_{f}]=5.4$ compared to that of the one-shot hit-and-run algorithm with $\mathbb{E}[\tau_{f}]=5.1$, in accordance with our intuition that larger range is beneficial.
The one-shot hit-and-run Algorithm 2 is $\sim 15$ times faster than the standard hit-and-run Algorithm 1 (with $10$ decorrelation steps) on an M1 CPU. Since the one-shot algorithm does not involve loops for each realization, it allows for efficient parallelization on GPUs. The standard hit-and-run algorithm, on the other hand, involves loops to execute the bisection, which prohibits effective parallelization on GPUs. As a result, the one-shot algorithm is $\sim 825$ faster than the standard algorithm on an A100 GPU. In applications we hence use the computationally more efficient one-shot hit-and-run Algorithm 2, given the almost equal performance in forecasting (cf.Figure 8).
@@
Figure 7. Empirical histogram of the effective range $\mathcal{R}$ of random feature maps when only good rows are used with $p_{g}=1$ for the standard hit-and-run Algorithm 1 (left) and the one-shot hit-and-run Algorithm 2 (right). Each histogram represents $10,000$ samples differing in the internal weights, the training data and the validation data. The feature dimension is $D_{r}=512$ and a regularization parameter of $\beta=2.79\times 10^{-5}$ is used with training data length $N=20,000$. For the standard hit-and-run Algorithm 1, $10$ decorrelation steps are used. For the standard hit-and-run algorithm we obtain a mean forecast time of $\mathbb{E}[\tau_{f}]=1.0$ with $\sigma[\tau_{f}]=0.03$. For the one-shot algorithm we obtain $\mathbb{E}[\tau_{f}]=0.42$ with $\sigma[\tau_{f}]=0.02$.
Figure 8. Empirical histogram of the forecast time $\tau_{f}$ when only good rows are used with $p_{g}=1$ for the standard hit-and-run Algorithm 1 and the one-shot hit-and-run Algorithm 2. The same random feature maps are used as in Figure 7. For the standard hit-and-run algorithm we obtain the mean forecast time $\mathbb{E}[\tau_{f}]=5.4$ with $\sigma[\tau_{f}]=1.5$. For the one-shot algorithm we obtain $\mathbb{E}[\tau_{f}]=5.1$ with $\sigma[\tau_{f}]=1.5$. The random feature maps shown here are the same as those in Figure 7.
5. Results for forecasting individual trajectories